An equipotential surface that surrounds a + 5.0 μC point charge has a radius of 5.0 cm. What is the potential of this surface? *




1
Expert's answer
2022-03-18T09:08:32-0400

The potential of electric field due point charge

"V=k\frac{q}{r}"

"V=9\times 10^9\times\frac{5.0\times 10^{-6}}{0.05}=9\times 10^5\:\rm V"